OpenStudy (llivingston2):

What is the equation in point-slope form of the line passing through (2, 5) and (−1, 8)? y+5=1(x+2) y+8=-1(x-1) y-5=-1(x-2) y-8=1(x-2)

OpenStudy (llivingston2):

@welshfella

OpenStudy (andres19):

use the formula m=y2-y1/x2-1 x1, y1 x2,y2 (2,5) (-1x,8)

OpenStudy (alivejeremy):

y+5=1(x+2)/.... not sure

OpenStudy (llivingston2):

how did you get that?

OpenStudy (andres19):

First, find the slope: = m=y2-y1/x2-x1 =8-5/2+1 3/3 =1 Use either point: (2,5) Remember the form: y - y1 = m ( x - x1) Substitute: y - 5 = 1 ( x - 2) y - 5 = (x -2) Ans.

OpenStudy (welshfella):

First find the slope (m) by perforing the calculation described by Andre

OpenStudy (llivingston2):

the slope is 1

OpenStudy (welshfella):

yea Andres has got it

OpenStudy (llivingston2):

oh alright

OpenStudy (llivingston2):

thank you andres19

OpenStudy (welshfella):

No hold one m = (8-5) / (-1-2) = 3/-3 = -1 slope is -1

OpenStudy (llivingston2):

I already turned it in and got it wrong

OpenStudy (welshfella):

oh sorry correct was y - 5 = -1(x - 2) do you understand how it was done. Thats the main thing.

OpenStudy (llivingston2):

yes i do now, andres had 3/3 instead of 3/-3

OpenStudy (welshfella):

yea (x1,y1) is (2,5) and (x2,y2) is (-1,8) so the slope is (y2-y1) / (x2-x1) = (8-5) / (-1-2) = 3/-3

OpenStudy (llivingston2):

can you tell me if this is right? Wendy played her favorite video game for 8 hours last week. Today, Wendy's parents restricted her to 6 additional hours each week for the next 9 weeks. Create the function, f(x), that models Wendy's total video game time, and explain what each number in the situation represents graphically, using complete sentences. i got : f(x) = 6x+8 f(x) = # hours x = #'d weeks

OpenStudy (andres19):

Welshfella he is right and sorry i forget the the negative

OpenStudy (llivingston2):

its ok

OpenStudy (welshfella):

f(x) is correct